Given csc(thada)=-5 and tan(thada)< 0 find the exact value for all reaming trig functions and then find the an thada to the nearest tenth of a degree with all restrictions that thada id a + angle
1
Expert's answer
2012-10-09T11:33:45-0400
sin=1/csc=-1/5 as tan<0 then cos>0
cos=sqrt(1-sin^2)=2sqrt(6)/5
tan=-1/2sqrt(6)
ctan=-2sqrt(6)
sec=5/2sqrt(6)